INAF queue observing, 2013 Nov03-Nov04

Observer: M. Fumana, R. Speziali
Telescope Operator: G. Becchetti
Support Astronomer: H. Khun (on call)
Instrument: (LBC / MODS)

Summary

At the sunset heavy clouds in the sky.
Humidity > 98%.
Fauked to observe with mods (one quadrant of red channel is down) and LBC (blue not collimating)

Data Summary

Project_RUN Objects Filter ExpTime Notes

Details

08:16 Humidity drops down.
  • The TO opens, to cool down the dome.

08:43 acqMODS mods.macsj0717_mos.acq

08:50 Slit and sky image have the 1st quadrant with not signal mods1r.0001-0002. Redo the acquisition

08:52 acqMODS mods.macsj0717_mos.acq
  • again mods1r.0003-0004 not valid 1st quadrant

09:01 acqMODS mods.macsj0717_mos.acq
  • again mods1r.0005-0006 not valid 1st quadrant

09:09 Switch to LBC

09:30 Run 2013B_6_A_copoint.ob

09:42 Communication problem: no images for dofpia

09:43 Run dofpia again

09:49 Communication problem: no images in /newdata available for dofpia
  • We call the SA
  • dofpia,/backout
  • exit
  • Bug in the new version of dofpia? The SA re-installed the previous version

10:03 Turn on the camera and connect it to LBT

10:04 Run 2013B_6_A_copoint.ob
  • The OB images are in /newdata
  • try dofpia now

10:07 dofpia,/first,/X2
  • dofpia runs but doesn't converge, defocus to large
  • TO clears adaptive optics
  • dofpia,/first,/X2,/blueonly
  • Blue doesn't converge. Z11+=1000 by hand

10:22 dofpia,/first,/X2,/blueonly
  • Huge Mirror temperature gradient and bad seeing are the reasons why the blue does not converge.
  • Humidity is raising again. A storm is coming

10:33 TO is closing the dome

-- MarcoFumana - 04 Nov 2013
Topic revision: r2 - 05 Nov 2013, RobertoSpeziali
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback